Find the term independent of x in the expansion (4x3 +1/2 x) raised to the power eight


1
Expert's answer
2019-10-31T07:01:46-0400

Let given binomial expansion be "(a+b)^n"

General term ,"T_{r+1}=\ ^{n}C_ra^{n-r}b^r"


"=\ ^{8}C_r(4x^3)^{8-r}({\frac{1}{2x}})^r"

Putting power of "x=0"

We get,


"24-4r=0\\or\\r=6"



So term will be

"T_{r+1}=T_7=\ ^8C_6(4^2)(\frac{1}{2})^6=\frac{28\times16}{64}"



"=7"